A 40 kg sled on frictionless ice is at rest next to you. You are standing on the ground. You push on it with a force of 10 N for 3 seconds. The sled moves away. You remain stationary.

a. How much total impulse was applied to the sled? (in N•s)

b. What is the final velocity of the sled?

c. How much impulse did the sled apply to you?

d. What is your final velocity?
Answer #1

(a). impulse=F.t =10×3 =30 N.s

(b) f.t=m(V-0)

30= 40×V. V=30/40 = 0.75m/s

V=0.75m/s

(c) impulse on me is zero because I was remain rest finally and initial was rest

(d) my velocity is zero because said that I remain stationary so final velocity is zero

If you are satisfied from my solution then plz like it Thank you